### v1

# Motion of a charged test particle around a static black hole in a monopole magnetic field

Abstract
We study the motion of a charged test particle in the spacetime with a spherically symmetric black hole which is immersed in a monopole magnetic field. We show that the radial motion of the charged test particle is governed by completely the same equation as that in the case of no magnetic field. This result implies that the black hole will acquire the electric charge if it is surrounded by the collisionless plasma composed of protons and electrons which obey the Maxwell velocity distribution. The drastically different situation appears in the tangential motions of charged test particles due to the magnetic field. The trajectory of a charged test particle in the black hole with the magnetic field of the order of 10 Gauss near the black hole, is confined on a very thin cone as long as the specific angular momentum of the particle is not much larger than the gravitational radius of the black hole times the speed of light. This result leads to a possibility that a plasma lump can hover over the black hole and is very hot, in the monopole magnetic field.
preprint: OCU-PHYS-618preprint: AP-GR-206
I Introduction
A black hole is defined as a complement of the causal past of the future null infinityHawking , and is not observable by its definitionNakao , whereas general relativity predicts that black holes exist in our universe. Black holes are predicted to be formed by the gravitational collapse of massive objects, and their mass may increase by absorbing surrounding matter. However, all we can observe is the black hole candidate that will form the black hole after infinite amount of time. Observable prediction related to the black hole, which is made by general relativity, is that the spacetime geometry around a stationary black hole is described by the Kerr-Newman family. A related issue is the black hole shadowLuminet . The Event Horizon Telescope (EHT) Collaboration reported that the black hole shadows at the center of the M87 galaxyEHT-M87-1 and the Milky Way galaxyEHT-MW-1 were photographed, but the debate over the reports by EHT Collaboration remain unresolvedMKM-1 ; MKM-2 ; MKM-3 .
Active Galactic Nuclei (AGN) are considered to be deeply related to black holes. The AGN are phenomena in each of which a narrow region at the center of a galaxy emits strong electromagnetic radiation that is almost equal to the entire galactic emission. The source of the radiative energy is thought to be the gravitational energy released by mass accretion onto the supermassive black hole Brandford-etal ; Lynden-Bell ; Bardeen ; Rees . Very long relativistic jets of plasma from AGNs are observed and their energy sources are regarded as one of important issues of the black hole astrophysics. Long jets implies that they should be steady for long time which is equal to or longer than the length of the jet divided by the speed of light. The Brandford-Znajek processBZ-process in which the rotational energy of the black hole is released through the neighboring magnetic field is regarded as a promising candidate of the energy source of the relativistic jet.
When we study physical phenomena around a black hole, the black hole is usually assumed to be electrically neutral. Wald, however, pointed out the possibility that in the case that a rotating black hole is immersed in a uniform magnetic field, since the rotation of the black hole causes a difference of an electrostatic potential between the event horizon and infinity, the black hole can acquire the electric charge, $8\pi\varepsilon_{0}G\vec{J}\cdot\vec{B}/c^{2}$, where $c$, $\varepsilon_{0}$, $G$, $\vec{J}$ and $\vec{B}$ are the speed of light, the electric permittivity of vacuum, Newton’s gravitational constant, the angular momentum of the black hole and the magnetic field, respectivelyWald-solution . Recently, King and Pringle stated that if the rotating black hole gets the electric charge of Wald’s value, the Brandford-Znajek process cannot provide energy sources of the steady relativistic jets from AGN’sKing-Pringle . Komissarov refuted their statementm and this issue is still an open question. Anyway, the electrification of black holes will be important in the consideration of the black hole astrophysicsKomissarov .
Zajaček et al. showed that even in the case of a non-rotating black hole with no magnetic field, the black hole can acquire an electric charge less than about $2\pi\varepsilon_{0}GMm_{\rm p}/e$ or two times of it, where $m_{\rm p}$ and $e$ are the mass of a proton and the elementary charge, respectivelyZajacek-etal . Since the analyses of Zajaček et al are basically based on the non-relativistic analysis, four of the present authors, KN, KM, HY, HI, studied the electrification of a static black hole by investigating the difference of the probabilities of a proton and an electron, whose initial velocity obey the Maxwell distribution, falling into the black holepaper-I . Hereafter we call Ref. paper-I NMYI paper. In NMYI paper, the maximal electric charge acquired by a static black hole is the same as that obtained by Zajaček et al, i.e., $2\pi\varepsilon_{0}GMm_{\rm p}/e$, and the radii of the Innermost Stable Circular Orbits (ISCO) of a proton and an electron can be very larger than that of a neutral test particle around a static charged black hole.
By those reasons mentioned above, in order to understand the physical situation around a black hole, it is crucially important to understand the motion of charged particles around the black hole immersed in electromagnetic field. In this paper, we study the motion of a charged test particle around a static black hole immersed in a monopole magnetic field. By using results obtained through this study, we investigate a possibility of the electrification of the black hole, and also make a comment on the temperature of the collisionless plasma surrounding the black hole. Here we should mention that Khan and Chen studied a charged test particle in the spacetime with a rotating black hole which is immersed in a split magnetic monopole field Khan-Chen . However, their interest is on the magnetic Penrose process, and very different from our present purposes.
This paper is organized as follows. In Sec. II, we briefly review the spherically symmetric black hole in monopole electromagnetic field. In Sec. III, we study the motion of a charged test particle around the spherically symmetric black hole in the monopole electromagnetic field. In Sec. IV, we show that a non-rotating black hole in a monopole magnetic field will acquire the positive electric charge as in the case without the magnetic field. In Sec. V, consider the effects of magnetic field on the temperature of the plasma surrounding the black hole by using the result obtained in Secs. II and III. Section VI is devoted for summary and discussion. In Appendix A, the probability of the test particle initially moving outward falling into the black hole is estimated. In Appendix B, we estimate the electric current on the surface of a lump of charged test particles.
In this paper, we adopt the abstract index notation in which the Latin indices, except for $t$ and $r$ which are assigned to the time and radial coordinates, represent a type of a tensor, whereas the Greek indices, except for $\theta$ and $\varphi$ which are assigned to the coordinates on a sphere, represent the components of a tensor with respect to the coordinate bases, and follow the sign convention of the metric tensor in Ref. Wald and adopt the SI unit. The covariant derivative and the ordinary derivative are denoted by $\nabla_{a}$ and $\partial_{a}$, respectively. The speed of light and Newton’s gravitational constant are denoted by $c$ and $G$, respectively.
II Static black hole in a monopole electromagnetic field
In this paper, we consider the electrification of a static black hole immersed in a monopole magnetic field. If there is no matter around the black hole, the spacetime geometry is described by the dyonic Reissner-Nordström (RN) solution whose infinitesimal world interval is given by
$ds^{2}=g_{\alpha\beta}dx^{\alpha}dx^{\beta}=-c^{2}f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right),$ (1)
with
$f(r)=1-\frac{2GM}{c^{2}r}+\frac{Q^{2}}{r^{2}}$ (2)
where $Q$ and $M$ are the charge parameter and gravitational mass in this system, respectively,
In the case that the electrified black hole is in a monopole magnetic field, the four-vector potential is given by
$A_{\mu}=-\frac{Q_{\rm E}}{4\pi\varepsilon_{0}r}\delta^{t}_{\mu}-\frac{Q_{\rm M}}{4\pi\varepsilon_{0}c}\cos\theta\delta^{\varphi}_{\mu},$ (3)
where $Q_{\rm E}$ and $Q_{\rm M}$ are the electric charge and the magnetic charge, of the black hole, respectively.
The charge parameter $Q$ of the RN solution is related to the electric charge $Q_{\rm E}$ and the magnetic charge $Q_{\rm M}$ in the manner,
$Q^{2}=\frac{G}{4\pi\varepsilon_{0}c^{4}}\left(Q_{\rm E}^{2}+Q_{\rm M}^{2}\right).$ (4)
In this paper, we focus on situations similar to the neighborhood of, for example, SgrA*, or the center of M87. The mass of SgrA* is equal to about $4\times 10^{6}M_{\odot}$ and the magnetic field in its neighborhood is estimated to be about 1010G EHT-MW-5 , whereas the mass of the black hole candidate in M87 is equal to about $6.5\times 10^{9}M_{\odot}$ and the magnetic field in its neighborhood is estimated be also about 1010G EHT-M87-5 . The magnetic charge $Q_{\rm M}$ which generates the magnetic field $B^{a}$ of the value 1010G at $r=Nr_{\rm g}$ is estimated as follows, where $N$ is a free parameter larger than one, and
$r_{\rm g}:=\frac{2GM}{c^{2}}$ (5)
is the gravitational radius of the black hole. We have
$B^{\alpha}=\frac{1}{2}\sqrt{c^{2}f(r)}\varepsilon^{tr\mu\nu}F_{\mu\nu}\delta^{\alpha}_{r}=\frac{\sqrt{f(r)}}{r^{2}\sin\theta}\left(\partial_{\theta}A_{\varphi}-\partial_{\varphi}A_{\theta}\right)\delta^{\alpha}_{r}=\frac{Q_{\rm M}\sqrt{f(r)}}{4\pi\varepsilon_{0}cr^{2}}\delta^{\alpha}_{r},$ (6)
where $\varepsilon^{abcd}$ is the skew tensor whose component is given as $\varepsilon^{0123}=1/\sqrt{-g}$ with $g$ being the determinant of the metric tensor. Hence, we obtain
$\frac{Q_{\rm M}}{4\pi\varepsilon_{0}c}=\frac{r^{2}B^{r}}{\sqrt{f(r)}}=\left(\frac{N^{5}}{N-1}\right)^{1\over 2}~r_{\rm g}^{2}B_{N},$ (7)
where
$B_{N}:=B^{r}|_{r=Nr_{\rm g}}.$ (8)
By using Eqs. (4) and (7), in the case of $Q_{\rm E}=0$, we have
$\frac{1}{r_{\rm g}}\times\frac{Q^{2}}{r}$ $=\frac{GQ_{\rm M}^{2}}{4\pi\varepsilon_{0}c^{4}r_{\rm g}^{2}}\left(\frac{r_{\rm g}}{r}\right)=\frac{4\pi\varepsilon_{0}Gr_{\rm g}^{2}}{c^{2}}\left(\frac{N^{5}}{N-1}\right)\left(\frac{r_{\rm g}}{r}\right)B_{N}^{2}$
$=1.1\times 10^{-23}\left(\frac{r_{\rm g}}{r}\right)\left(\frac{N^{5}}{N-1}\right)\left(\frac{B_{N}}{10^{-3}{\rm T}}\right)^{2}\left(\frac{M}{4\times 10^{6}M_{\odot}}\right)^{2}.$ (9)
As will be mentioned later, the energy of the electric field generated by the electric charge acquired by the black hole is also very small. Hence the spacetime geometry is well approximated by the Schwarzschild solution whose infinitesimal world interval is given by Eq. (1) with not Eq. (2) but
$f(r)=1-\frac{r_{\rm g}}{r}.$ (10)
Thus, hereafter, we assume the Schwarzschild metric.
III Motion of a charged test particle in monopole field
In this section, we study the motion of a charged test particle in the non-rotating black hole spacetime with the four-vector potential Eq. (3). The world line of a test particle is given in the form $x^{\alpha}=x^{\alpha}(\tau)$, where $\tau$ is the proper time of the test particle, and then the components of the four-velocity $u^{a}$ of the test particle with respect to the coordinate basis is defined as
$u^{\alpha}=\frac{dx^{\alpha}}{d\tau}.$ (11)
The four-velocity $u^{a}$ satisfies the normalization condition
$u^{a}u_{a}=-c^{2}.$ (12)
The equation of motion for a charged test particle with the mass $m$ and the electric charge $q$ is
$mu^{b}\nabla_{b}u_{a}=qF_{ab}u^{b},$ (13)
where $\nabla_{b}$ and is the covariant derivative, whereas $F_{ab}$ is the electromagnetic field strength tensor defined as
$F_{ab}=\nabla_{a}A_{b}-\nabla_{b}A_{a}=\partial_{a}A_{b}-\partial_{b}A_{a},$ (14)
where $\partial_{b}$ is the ordinary derivative.
III.1 Conserved quantities
If a Killing vector, $\xi^{a}$, exists, we have
$u^{b}\nabla_{b}\left(mu_{a}\xi^{a}\right)$ $=\left(mu^{b}\nabla_{b}u_{a}\right)\xi^{a}+mu^{a}u^{b}\nabla_{b}\xi_{a}$
$=q\left(\nabla_{a}A_{b}-\nabla_{b}A_{a}\right)\xi^{a}u^{b}$
$=q\left[\nabla_{a}\left(A_{b}+\nabla_{b}\chi\right)-\nabla_{b}\left(A_{a}+\nabla_{a}\chi\right)\right]\xi^{a}u^{b}$
$=q\left\{{\cal L}_{\xi}\left(A_{b}+\nabla_{b}\chi\right)-\nabla_{b}\left[\left(A_{a}+\nabla_{a}\chi\right)\xi^{a}\right]\right\}u^{b},$ (15)
where we have used Eq. (13) and the Killing equation $\nabla_{(a}\xi_{b)}=0$ in the second equality. Thus we have
$u^{b}\nabla_{b}\Bigl\{\bigl[mu_{a}+q\left(A_{a}+\nabla_{a}\chi\right)\bigr]\xi^{a}\Bigr\}=q\left[{\cal L}_{\xi}\left(A_{a}+\nabla_{a}\chi\right)\right]u^{a}.$ (16)
If ${\cal L}_{\xi}\left(A_{b}+\nabla_{b}\chi\right)=0$ holds, this equation implies that $\bigl[mu_{a}+q\left(A_{a}+\nabla_{a}\chi\right)\bigr]\xi^{a}$ is conserved.
In the spacetime whose metric is given by Eq. (1), the time coordinate basis vector $(\partial/\partial t)^{a}$ is a Killing vector, and there is the $SO(3)$ isometry whose generators are three Killing vectors given below. By defining the Cartesian coordinates $(x,y,z)$ for three dimensional space as
$(x,y,z)=(r\sin\theta\cos\varphi,r\sin\theta\sin\varphi,r\cos\theta),$ (17)
the three Killing vectors are given as
${L}_{(x)}^{a}$ $=y\left(\frac{\partial}{\partial z}\right)^{a}-z\left(\frac{\partial}{\partial y}\right)^{a}=-\sin\varphi\left(\frac{\partial}{\partial\theta}\right)^{a}-\cot\theta\cos\varphi\left(\frac{\partial}{\partial\varphi}\right)^{a},$ (18)
${L}_{(y)}^{a}$ $=z\left(\frac{\partial}{\partial x}\right)^{a}-x\left(\frac{\partial}{\partial z}\right)^{a}=\cos\varphi\left(\frac{\partial}{\partial\theta}\right)^{a}-\cot\theta\sin\varphi\left(\frac{\partial}{\partial\varphi}\right)^{a},$ (19)
${L}_{(z)}^{a}$ $=x\left(\frac{\partial}{\partial y}\right)^{a}-y\left(\frac{\partial}{\partial x}\right)^{a}=\left(\frac{\partial}{\partial\varphi}\right)^{a},$ (20)
where $\left(\dfrac{\partial}{\partial x^{\mu}}\right)^{a}$ denotes a coordinate basis vector. These Killing vectors generate the rotations around $x$-axis, $y$-axis and $z$-axis, respectively.
We introduce two functions, $\chi_{(x)}$ and $\chi_{(y)}$, defined as solutions of the following differential equations.
$L_{(x)}^{a}\nabla_{a}\chi_{(x)}$ $=-\frac{Q_{\rm M}}{\sin\theta}(1-\cos\theta)\cos\varphi,$ (21)
$L_{(y)}^{a}\nabla_{a}\chi_{(y)}$ $=-\frac{Q_{\rm M}}{\sin\theta}(1-\cos\theta)\sin\varphi$ (22)
The derivatives in these equations are the ordinary differentiations along the integral curves of the Killing vectors $L_{(x)}^{a}$ and $L_{(y)}^{a}$, respectively, and hence the solutions exist.
Then the following equations hold.
${\cal L}_{\partial\over\partial t}A_{a}$ $=0,$ (23)
${\cal L}_{L_{(x)}}\left(A_{a}+\nabla_{a}\chi_{(x)}\right)$ $=0,$ (24)
${\cal L}_{L_{(y)}}\left(A_{a}+\nabla_{a}\chi_{(y)}\right)$ $=0,$ (25)
${\cal L}_{L_{(z)}}A_{a}$ $=0,$ (26)
where note that both of ${\cal L}_{L_{(x)}}A_{a}$ and ${\cal L}_{L_{(y)}}A_{a}$ do not vanish. Thus we have the four conserved quantities given as
$E$ $:=-\left(u_{a}+\frac{q}{m}A_{a}\right)\left(\frac{\partial}{\partial t}\right)^{a}$
$=-u_{t}+\frac{qQ_{\rm E}}{4\pi\varepsilon_{0}mcr},$ (27)
$L_{(x)}$ $:=\left[u_{a}+\frac{q}{m}\left(A_{a}+\nabla_{a}\chi_{(x)}\right)\right]L_{(x)}^{a}$
$~=-u_{\theta}\sin\varphi-u_{\varphi}\cot\theta\cos\varphi-\frac{qQ_{\rm M}}{4\pi\varepsilon_{0}mc}\sin\theta\cos\varphi,$ (28)
$L_{(y)}$ $:=\left[u_{a}+\frac{q}{m}\left(A_{a}+\nabla_{a}\chi_{(y)}\right)\right]L_{(y)}^{a}$
$~=u_{\theta}\cos\varphi-u_{\varphi}\cot\theta\sin\varphi-\frac{qQ_{\rm M}}{4\pi\varepsilon_{0}mc}\sin\theta\sin\varphi,$ (29)
$L_{(z)}$ $:=\left(u_{a}+\frac{q}{m}A_{a}\right)L_{(z)}^{a}$
$=u_{\varphi}-\frac{qQ_{\rm M}}{4\pi\varepsilon_{0}mc}\cos\theta$ (30)
where $E$, $L_{(x)}$, $L_{(y)}$ and $L_{(z)}$ correspond to the specific energy, the $x$-, $y$- and $z$-components of the specific angular momentum, respectively.
From Eqs. (28)–(30), we have
$u_{\theta}^{2}+\frac{u_{\varphi}^{2}}{\sin^{2}\theta}=L_{(x)}^{2}+L_{(y)}^{2}+L_{(z)}^{2}-\left(\frac{qQ_{\rm M}}{4\pi\varepsilon_{0}mc}\right)^{2}=:K^{2}.$ (31)
III.2 Motion in radial direction
For notational simplicity, we introduce the dimensionless quantities defined as
$R$ $:=\frac{r}{r_{\rm g}},$ (32)
${\cal T}$ $:=\frac{c\tau}{r_{\rm g}},$ (33)
${\cal E}$ $:=\frac{E}{c},$ (34)
$\left({\cal L}_{(x)},{\cal L}_{(y)},{\cal L}_{(z)}\right)$ $:=\left(\frac{L_{(x)}}{cr_{\rm g}},\frac{L_{(y)}}{cr_{\rm g}},\frac{L_{(z)}}{cr_{\rm g}}\right),$ (35)
${\cal K}$ $:=\frac{K}{cr_{\rm g}},$ (36)
${\cal Q}_{\rm E}$ $:=\frac{qQ_{\rm E}}{4\pi\varepsilon_{0}mc^{2}r_{\rm g}},$ (37)
${\cal Q}_{\rm M}$ $:=\frac{qQ_{\rm M}}{4\pi\varepsilon_{0}mc^{2}r_{\rm g}},$ (38)
From the normalization of the four velocity, Eq. (12), we have
$\left(\frac{dR}{d{\cal T}}\right)^{2}+U(R;{\cal E},{\cal K}^{2},{\cal Q}_{\rm E})=0,$ (39)
where
$U(R;{\cal E},{\cal K}^{2},{\cal Q}_{\rm E}):=-\left({\cal E}-\frac{{\cal Q}_{\rm E}}{R}\right)^{2}+\left(1-\frac{1}{R}\right)\left(1+\frac{{\cal K}^{2}}{R^{2}}\right).$ (40)
Equation (39) is identical to the corresponding equation for the charged test particle in the Schwarzschild spacetime with no magnetic charge $Q_{\rm M}=0$ but with non-vanishing electric charge $Q_{\rm E}\neq 0$. No effect of the monopole magnetic field appears in the radial motion of the charged test particle, and hence, for example, the radius of ISCO is unchanged from that with no magnetic field. [^2] It should be noted that ISCO radius the rotating black hole immersed in the asymptotically homogeneous magnetic field can be very different from the case without magnetic field IHK2012 .
We write $U(R;{\cal E},{\cal K}^{2},{\cal Q}_{\rm E})$ in the form
$U(R;{\cal E},{\cal K}^{2},{\cal Q}_{\rm E})=-\left[{\cal E}-\frac{{\cal Q}_{\rm E}}{R}+\sqrt{\left(1-\frac{1}{R}\right)\left(1+\frac{{\cal K}^{2}}{R^{2}}\right)}\right]\left[{\cal E}-V(R;{\cal K}^{2},{\cal Q}_{\rm E})\right]$ (41)
where
$V(R;{\cal K}^{2},{\cal Q}_{\rm E})=\frac{{\cal Q}_{\rm E}}{R}+\sqrt{\left(1-\frac{1}{R}\right)\left(1+\frac{{\cal K}^{2}}{R^{2}}\right)}.$ (42)
By noting that
$f(r)\frac{dt}{d\tau}={\cal E}-\frac{{\cal Q}_{\rm E}}{R}>0~~~{\rm for}~r>r_{\rm g}$ (43)
must hold, we find that the allowed domain for the motion of the charged test particle is determined by the inequality
$V(R;{\cal K}^{2},{\cal Q}_{\rm E})\leq{\cal E}.$ (44)
III.3 Motion in angular directions
Here, we adopt the coordinates in which the angular momentum of the charged test particle is parallel to the $z$-axis, i.e., $L_{(x)}=L_{(y)}=0$. We have, from Eqs. (28) and (29),
$u_{\theta}=-L_{(x)}\sin\varphi+L_{(y)}\cos\varphi,$ (45)
and hence $u_{\theta}=r^{2}\dfrac{d\theta}{d\tau}=0$ holds in the coordinate we adopt. The motion of the charged test particle is restricted in a hypersurface of constant polar angle
$\theta=\vartheta.$ (46)
Substituting $L_{(x)}=L_{(y)}=0$, $u_{\theta}=0$ and $\theta=\vartheta$ into Eqs. (28) and (29), we have
$u_{\varphi}\cos\vartheta+{\cal Q}_{\rm M}cr_{\rm g}\sin^{2}\vartheta=0.$ (47)
Substituting this equation into Eq. (30) to eliminate $u_{\varphi}$, we have
${\cal L}_{(z)}\cos\vartheta+{\cal Q}_{\rm M}=0$ (48)
We find, from this equation, that the sign of ${\cal L}_{(z)}$ is opposite to that of ${\cal Q}_{\rm M}$ in $0\leq\vartheta<\pi/2$, whereas the sign of ${\cal L}_{(z)}$ is the same as that of ${\cal Q}_{\rm M}$ in $\pi/2\leq\vartheta\leq\pi$. Eliminating ${\cal Q}_{\rm M}$ from Eq. (47) by using Eq. (48), we obtain
$u_{\varphi}={\cal L}_{(z)}cr_{\rm g}\sin^{2}\vartheta.$ (49)
This equation implies that the sign of $\dfrac{d\varphi}{d\tau}$ is the same as that of ${\cal L}_{(z)}$. Equation (47) is rewritten in the form
$u_{\varphi}=-{\cal Q}_{\rm M}cr_{\rm g}\sin\vartheta\tan\vartheta.$ (50)
As expected, the direction of rotational motion of the charged test particle is determined by the sign of ${\cal Q}_{\rm M}$ which is proportional to the product of $q$ and $Q_{\rm M}$.
We have, from Eqs. (31), (36) and (48),
$\tan^{2}\vartheta=\frac{{\cal K}^{2}}{{\cal Q}_{\rm M}^{2}}.$ (51)
By using Eq. (7), the dimensionless magnetic charge parameter is estimated to be
${\cal Q}_{\rm M}=\frac{qQ_{\rm M}}{4\pi\varepsilon_{0}mc^{2}r_{\rm g}}=3.8\times 10^{6}\left(\frac{N^{5}}{N-1}\right)^{1\over 2}\left(\frac{q}{e}\right)\left(\frac{m_{\rm p}}{m}\right)\left(\frac{M}{4\times 10^{6}M_{\odot}}\right)\left(\frac{B_{N}}{10^{-3}{\rm T}}\right)$ (52)
where $e$ and $m_{\rm p}$ are the elementary charge and the mass of a proton, respectively. In the neighborhood of SgrA*, ${\cal Q}_{\rm M}$ of both the proton and the electron is very large. Thus, we find, from Eq. (52), that, in the neighborhood of SgrA* and in the neighborhood of the center in M87, $|\vartheta|\ll 1$ holds for the charged test particle of ${\cal K}^{2}$ which is much less than ${\cal Q}_{\rm M}^{2}$.
IV Electric charge acquired by the black hole in monopole magnetic field
In this section, we consider the electrification of a static black hole in the monopole magnetic field, ${\cal Q}_{\rm M}\neq 0$. It is determined by the effective potential $V(R;{\cal E},{\cal K}^{2},{\cal Q}_{\rm E})$ defined as Eq. (42) whether the charged test particle falls into the black hole. Since the effective potential $V(R;{\cal E},{\cal K}^{2},{\cal Q}_{\rm E})$ is completely the same as that in the case without the magnetic field studied in NMYI paperpaper-I , the results on the electrification of a static black hole obtained in NMYI paper is also obtained in the present case.
We assume charged test particles compose plasma which is in local thermal equilibrium, and the velocity of a charged test particle is given by the Maxwell distribution
$F(\bm{v})$ $=\left(\frac{m}{2\pi k_{\rm B}T}\right)^{3\over 2}\exp\left(-\frac{m}{2k_{\rm B}T}\bm{v}^{2}\right),$ (53)
where $k_{\rm B}$ and $T$ are the Boltzmann constant and the temperature, whereas $\bm{v}$ is the spatial components of the four velocity $u^{a}$ with respect to the orthonormal basis associated to the static observer
$e_{(t)\mu}$ $=\left(-\sqrt{f(r)},0,0,0\right),$ (54)
$e_{(r)\mu}$ $=\left(0,\frac{1}{\sqrt{f(r)}},0,0\right),$ (55)
$e_{(\theta)\mu}$ $=\left(0,0,r,0\right),$ (56)
$e_{(\varphi)\mu}$ $=\left(0,0,0,r\sin\theta\right).$ (57)
Hence, we have
$v_{(r)}$ $=e_{(r)\mu}u^{\mu}=\frac{1}{\sqrt{f(r)}}\frac{dr}{d\tau},$ (58)
$v_{(\theta)}$ $=e_{(\theta)\mu}u^{\mu}=r\frac{d\theta}{d\tau}=\frac{u_{\theta}}{r},$ (59)
$v_{(\varphi)}$ $=e_{(\varphi)\mu}u^{\mu}=r\sin\theta\frac{d\varphi}{d\tau}=\frac{u_{\varphi}}{r\sin\theta}.$ (60)
From Eq. (31), the following equality holds,
$v_{(\theta)}^{2}+v_{(\varphi)}^{2}=\frac{c^{2}{\cal K}^{2}}{R^{2}}.$ (61)
Note that the assumption of the Maxwell distribution is justified only if $k_{\rm B}T\ll mc^{2}$ holds.
We are interested in the electrification of an electrically neutral black hole but in a monopole magnetic field. Thus, first we consider the case with ${\cal Q}_{\rm E}=0$ and ${\cal Q}_{\rm M}\neq 0$. In this case, the effective potential for the charged test particle, which is defined as Eq. (42) becomes
$V(R;{\cal K}^{2},{\cal Q}_{\rm E}=0)=\sqrt{\left(1-\frac{1}{R}\right)\left(1+\frac{{\cal K}^{2}}{R^{2}}\right)}.$ (62)
The effective potential $V$ takes extrema at the roots of the equation
$\dfrac{\partial V(R;{\cal K}^{2},{\cal Q}_{\rm E}=0)}{\partial R}=0,$ (63)
or equivalently
$R^{2}-2{\cal K}^{2}R+3{\cal K}^{2}=0.$ (64)
The roots of this equation are given as
$R=R_{\pm}:={\cal K}^{2}\left(1\pm\sqrt{1-\frac{3}{{\cal K}^{2}}}\right).$ (65)
We find from this equation that, in the case of ${\cal K}^{2}<3$, there is no extremum of $V(R;{\cal K}^{2},{\cal Q}_{\rm E}=0)$, and hence the test particles with ingoing initial velocity necessarily enters the black hole.
In the case of ${\cal K}^{2}\geq 3$, the sufficient condition for the test particle to fall into the black hole is $v_{(r)}<0$ and ${\cal E}>V(R_{-};{\cal K}^{2},{\cal Q}_{\rm E}=0)$. The normalization condition for the four-velocity, Eq. (12), is rewritten in the form,
$\frac{v_{(r)}^{2}}{c^{2}}=\frac{{\cal E}^{2}}{1-1/R}-1-\frac{{\cal K}^{2}}{R^{2}}.$ (66)
Hence the condition ${\cal E}>V(R_{-};{\cal K}^{2},{\cal Q}_{\rm E}=0)$ is equivalent to
$\frac{v_{(r)}^{2}}{c^{2}}>\frac{V^{2}(R_{-};{\cal K}^{2},{\cal Q}_{\rm E}=0)}{1-1/R}-1-\frac{{\cal K}^{2}}{R^{2}},$ (67)
where, from Eq. (65), we have
$V^{2}(R_{-};{\cal K}^{2},{\cal Q}_{\rm E}=0)=1+\frac{{\cal K}^{2}}{27}\left[2-\frac{9}{{\cal K}^{2}}+2\left(1-\frac{3}{{\cal K}^{2}}\right)^{3\over 2}\right].$ (68)
Using this expression, Eq. (67) is rewritten as
$2{\cal K}^{2}\left(1-\frac{3}{{\cal K}^{2}}\right)^{3\over 2}<9+27\left(1-\frac{1}{R}\right)\frac{v_{(r)}^{2}}{c^{2}}-\frac{27}{R}-\left[2-\frac{27}{R^{2}}\left(1-\frac{1}{R}\right)\right]{\cal K}^{2}.$ (69)
Since the L.H.S. of this inequality is non-negative, the R.H.S. should be non-negative so that there exists ${\cal K}^{2}$ which satisfies this inequality. Since we focus on the case of $R\gg 1$, the following inequality should hold so that Eq. (67) holds.
${\cal K}^{2}\leq\frac{9}{2}\left[1+3\left(1-\frac{1}{R}\right)\frac{v_{(r)}^{2}}{c^{2}}-\frac{3}{R}\right]\left[1-\frac{27}{2R^{2}}\left(1-\frac{1}{R}\right)\right]^{-1}.$ (70)
Assuming Eq. (70), by taking the square of both sides of Eq. (69), a bit complicated manipulation leads
$c_{0}+c_{1}{\cal K}^{2}+c_{2}{\cal K}^{4}+c_{3}{\cal K}^{6}>0,$ (71)
where
$c_{0}$ $=4,$ (72)
$c_{1}$ $=-1+9\left[\left(1-\frac{1}{R}\right)\frac{v_{(r)}^{2}}{c^{2}}-\frac{1}{R}\right]\left\{2+3\left[\left(1-\frac{1}{R}\right)\frac{v_{(r)}^{2}}{c^{2}}-\frac{1}{R}\right]\right\},$ (73)
$c_{2}$ $=-4\left[\left(1-\frac{1}{R}\right)\frac{v_{(r)}^{2}}{c^{2}}-\frac{1}{R}\right]+\frac{18}{R^{2}}\left(1-\frac{1}{R}\right)\left\{1+3\left[\left(1-\frac{1}{R}\right)\frac{v_{(r)}^{2}}{c^{2}}-\frac{1}{R}\right]\right\},$ (74)
$c_{3}$ $=-\frac{1}{R^{2}}\left(1-\frac{1}{R}\right)\left[4-\frac{27}{R^{2}}\left(1-\frac{1}{R}\right)^{2}\right].$ (75)
By solving Eq. (71) with respect to ${\cal K}^{2}$, we obtain the following inequality
${\cal K}^{2}<{\cal K}_{\rm max}^{2}\left(v_{(r)}^{2},R\right),$ (76)
where, in the case that $v_{(r)}^{2}/c^{2}$ and $1/R$ are the same order as one another and much less than unity, [^3] We assumed $v_{(r)}^{2}/c^{2}\gg 1/R$ in NMYI paper. The assumption in NMYI paper is equivalent to assume that the test particles are gravitationally unbound. By contrast, in this paper, we do not assume so, although the main result Eq. (80) is the same as that in NMYI paper. we have
${\cal K}_{\rm max}^{2}\left(v_{(r)}^{2},R\right)=4+8\left(\frac{v_{(r)}^{2}}{c^{2}}-\frac{1}{R}\right)+{\cal O}\left(\frac{v_{(r)}^{4}}{c^{4}},\frac{1}{R^{2}}\right).$ (77)
Note that, in the case of $R\gg 1$ and $v^{2}_{(r)}/c^{2}\ll 1$, Eq. (70) holds, if Eq. (76) holds.
By using Eq. (61), we rewriter $v_{(\theta)}$ and $v_{(\varphi)}$ as
$v_{(\theta)}$ $=\frac{c{\cal K}}{R}\cos\phi,$ (78)
$v_{(\varphi)}$ $=\frac{c{\cal K}}{R}\sin\phi,$ (79)
where $0\leq\phi<2\pi$.
The test particle which initially moves outward will fall into the black hole only if it is bounded. As shown in Appendix A, the probability of a bounded particle initially moving outward for falling into the black hole is much less than that of a particle initially moving inward. Hence hereafter we consider only the probability $P$ of a particle initially moving inward for falling into the black hole. It is given as
$P$ $=\int_{-\infty}^{0}dv_{(r)}\left(\frac{m}{2\pi k_{\rm B}T}\right)^{1\over 2}\exp\left(-\frac{m}{2k_{\rm B}T}v_{(r)}^{2}\right)$
$~~\times\int_{0}^{2\pi}d\phi\int_{0}^{{\cal K}_{\rm max}^{2}\left(v_{(r)}^{2},R\right)}d{\cal K}^{2}\frac{mc^{2}}{4\pi k_{\rm B}TR^{2}}\exp\left(-\frac{mc^{2}}{2k_{\rm B}TR^{2}}{\cal K}^{2}\right)$
$=\int_{-\infty}^{0}dv_{(r)}\left(\frac{m}{2\pi k_{\rm B}T}\right)^{1\over 2}\exp\left(-\frac{m}{2k_{\rm B}T}v_{(r)}^{2}\right)\left\{1-\exp\left[-\frac{mc^{2}}{2k_{\rm B}TR^{2}}{\cal K}_{\rm max}^{2}\left(v_{(r)}^{2},R\right)\right]\right\}$
$\simeq\frac{mc^{2}}{k_{\rm B}TR^{2}},$ (80)
where we have used Eq. (77) in the last equality.
Hereafter the test particle is assumed to be a proton or an electron. Any quantity of a proton is denoted by the character with a subscript “p”, whereas that of an electron is denoted by the character with a subscript “e”. Then the excess of the probability of the proton to enter the black hole is given as
$\Delta P=P_{\rm p}-P_{\rm e}\simeq\frac{1}{R^{2}}\left(\frac{m_{\rm p}c^{2}}{k_{\rm B}T_{\rm p}}-\frac{m_{\rm e}c^{2}}{k_{\rm B}T_{\rm e}}\right).$ (81)
If $m_{\rm p}/T_{\rm p}>m_{\rm e}/T_{\rm e}$ holds, $\Delta P$ is positive and the black hole will be positively electrified. By contrast, if $m_{\rm p}/T_{\rm p}<m_{\rm e}/T_{\rm e}$ holds, $\Delta P$ is negative and the black hole will be negatively electrified. If $m_{\rm p}/T_{\rm p}=m_{\rm e}/T_{\rm e}$ holds, no electrification of the black hole will occur. If the temperature of the protons and that of the electrons are identical to each other, the black hole will be positively electrified because of $m_{\rm p}\gg m_{\rm e}$. This result is completely the same as that obtained in NMYI paper.
Although the detailed derivation is not shown in this paper, the maximal charge acquired by the black hole in the monopole magnetic field is also the same as that revealed in NMYI paper, i.e., ${\cal Q}_{\rm E}=1/2$ for a static black hole with no magnetic field. This value is not so large but the radius of ISCO of the charged test particle can be very different from that of the electrically neutral particle.
It is easy to see that even in the case that the magnetic field is a split monopole, the present result is unchanged, since there is no magnetic charge ${\cal Q}_{\rm M}$ in the effective potential Eq. (42).
V Collisionless plasma hovering on the black hole
For simplicity, in this section, we focus on the case of the electrically charge neutral black hole, ${\cal Q}_{\rm E}=0$, although the effect of the electric charge acquired by the black hole is an interesting issue. As shown in the subsection III.3, a charged test particle does not move on an equatorial plane but on a very thin cone.
V.1 “Temperature” of bounded collisionless plasma
Now, let us consider a lump of collisionless plasma bounded by the gravity of the central black hole, which is assumed to be far from the black hole,
$1\ll R_{\rm min}<R<R_{\rm max}.$ (82)
The energy density of the plasma is assumed to be much less than that of the magnetic field and hence too small to affect the spacetime geometry. Thus the charged particles in the plasma can be treated as test particles. In such a situation, the energy equation for a test particle, which is derived from the normalization condition of the four velocity, is approximately given as
$\left(\frac{dR}{d{\cal T}}\right)^{2}-\frac{1}{R}+\frac{{\cal K}^{2}}{R^{2}}=E,$ (83)
where $E={\cal E}^{2}-1$. Iff
$\frac{1}{4{\cal K}^{2}}\leq E<0$ (84)
holds, the motion of the test particle is bounded in the domain
$R_{-}\leq R\leq R_{+},$ (85)
where
$R_{\pm}=-\frac{1}{2E}\left(1\pm\sqrt{1+4E{\cal K}^{2}}\right).$ (86)
Then, from Eq. (82), the allowed $E$ and ${\cal K}^{2}$ are determined by
$R_{\rm min}\leq R_{-}\leq R_{+}\leq R_{\rm max}.$ (87)
From Eq. (84), the inequality, $E<0$, should hold. The condition $R_{\rm min}\leq R_{-}$ leads to two inequalities,
$|E|\leq\frac{1}{2R_{\rm min}}$ (88)
and
${\cal K}^{2}\geq-R_{\rm min}^{2}|E|+R_{\rm min},$ (89)
whereas the condition $R_{+}\leq R_{\rm max}$ leads to
$|E|\geq\frac{1}{2R_{\rm max}}$ (90)
and
${\cal K}^{2}\geq-R_{\rm max}^{2}|E|+R_{\rm max}.$ (91)
An inequality
${\cal K}^{2}\leq\frac{1}{4|E|}$ (92)
is derived from Eq. (84). Inequalities (88)–(92) determine the allowed $|E|$ and ${\cal K}^{2}$ (see a schematic diagram in Fig. 1).
Figure 1: A schematic diagram. Equation (87) is satisfied in the shaded region.
In the situation we consider, the square of the dimensionless velocity, ${\cal V}^{2}$, is given by
${\cal V}^{2}=\left(\frac{dR}{d{\cal T}}\right)^{2}+\frac{{\cal K}^{2}}{R^{2}}.$ (93)
Since the motion of the test particle is bounded, the time average of $(dR/d{\cal T})^{2}$ is given as
$\left\langle\left(\frac{dR}{d{\cal T}}\right)^{2}\right\rangle$ $=\lim_{T\rightarrow\infty}\frac{1}{2T}\int_{-T}^{T}\left(\frac{dR}{d{\cal T}}\right)^{2}d{\cal T}$
$=\lim_{T\rightarrow\infty}\frac{1}{2T}\int_{-T}^{T}\left[\frac{d}{d{\cal T}}\left(R\frac{dR}{d{\cal T}}\right)-R\frac{d^{2}R}{d{\cal T}^{2}}\right]d{\cal T}$
$=\lim_{T\rightarrow\infty}\frac{1}{2T}\int_{-T}^{T}\left(\frac{1}{2R}-\frac{{\cal K}^{2}}{R^{2}}\right)d{\cal T}$
$=\left\langle\frac{1}{2R}\right\rangle-\left\langle\frac{{\cal K}^{2}}{R^{2}}\right\rangle,$ (94)
where, in the third equality, we have used the equation of motion derived by differentiating Eq. (83) with respect to $\cal T$},
$\frac{d^{2}R}{d{\cal T}^{2}}=-\frac{1}{2R^{2}}+\frac{{\cal K}^{2}}{R^{3}}.$ (95)
From Eqs. (93) and (94), we have
$\langle{\cal V}^{2}\rangle=\left\langle\frac{1}{2R}\right\rangle,$ (96)
or equivalently,
$\langle{\cal V}^{2}\rangle=-E.$ (97)
From the condition Eq. (87), we obtain
$\frac{1}{2R_{\rm max}}\leq\langle{\cal V}^{2}\rangle\leq\frac{1}{2R_{\rm min}}.$ (98)
This inequality implies that if the plasma is confined in the domain $R_{\rm min}\leq R\leq R_{\rm max}$, the time average of its kinetic energy is restricted by
$\frac{mc^{2}}{4R_{\rm max}}\leq\left\langle\frac{1}{2}mc^{2}{\cal V}^{2}\right\rangle\leq\frac{mc^{2}}{4R_{\rm min}}.$ (99)
In the case of the proton, the time average of the kinetic energy is bounded as
$2.35\times 10^{6}\left(\frac{100}{R_{\rm max}}\right){\rm eV}\leq\left\langle\frac{1}{2}m_{\rm p}c^{2}{\cal V}^{2}\right\rangle\leq 2.35\times 10^{7}\left(\frac{10}{R_{\rm min}}\right){\rm eV}.$ (100)
The velocities of protons and electrons on the same orbit are equal to each other since they are governed by the gravity of the black hole. Thus the average of the kinetic energy of protons is about 1833 times larger than that of an electron.
Let us consider a lump of collisionless plasma composed of protons and electrons, whose size is assumed to be, for example, the order of the gravitational radius $r_{\rm g}$. Its radial extent, $R_{\rm max}-R_{\rm min}$, is of order one and the tangential extent is $2R_{\rm min}\Theta=1$. Here note that the half apex angle $\vartheta$ of the cone on which the motion of the charged test particle is restricted is much less $\Theta$ in the situation of our interest, $10\lesssim R_{\rm min}<R_{\rm max}\lesssim 100$ (see Eqs. (51) and (52)). This means that the Larmor radius of the charged test particle is much less than the size of the plasma lump, since the Larmor radius is less than $R_{\rm max}\sin\vartheta$. Hence the random motions of the charged test particles in the plasma lump are Larmor one if the plasma is really collisionless. In this case, the time average of the kinetic energy estimated in Eq. (100) might be regarded as the “temperature” of the plasma lump. Here it is very remarkable fact that the kinetic energy of a charged test particle is equal to that of the “Keplerian” motion, since the energy equation of the charged test particle is given by Eq. (83). Hence, the temperature of the plasma lump can be very high in the neighborhood of the black hole in the monopole magnetic field. The origin of this temperature is the gravity.
As mentioned, in the case that the plasma is composed of protons and electrons, their kinetic energies should be very different from each other due to the mass difference. Even if the velocity distribution of the protons and that of the electrons becomes Maxwell one with the same temperature by the thermalization through some plasma instability, the domain allowed for the motions of protons and those of electrons will be very different. Hence, the local thermal equilibrium seems to be unachievable in the plasma lump hovering on the black hole.
It should be noted that the average velocity of electrons will be equal to that of protons in a plasma lump. This situation is very different from the assumed situation in the discussion of the electrification of a static black hole, where we have assumed that the velocity distribution of electrons and that of protons are Maxwell one with not so different temperatures from each other. Hence, we should note that lumps of collisionless plasma will not be suppliers of electric charge to the black hole.
Here note that the global monopole magnetic field is not a necessary condition for the hovering motion. If the configuration of the magnetic field only in the domain of the motion of the charged test particle is locally similar to the monopole magnetic field, the charged test particle will hover on the black hole. In the stationary axi-symmetric magnetosphere, the neighborhood of the rotational symmetry axis of a rotating black hole may be the case (see for example, Fig. 3 in Ref. MG2004 ). Thus, in the realistic situation like as the neighborhood of SgrA*, the heating of a plasma lump may occur in the vicinity of the symmetry axis of rotating black hole.
V.2 Collisionless condition
The two-body relaxation time for the plasma composed of protons and electrons is estimated as the time scale for which the accumulated changes in velocity through multiple two-body Coulomb scatterings becomes the same order of the initial velocity. The shortest time scale is given by the electron-electron scattering.
$T_{\rm relax}=\frac{4\pi\varepsilon_{0}^{2}m_{\rm e}^{2}\ell^{3}v_{\rm e}^{3}}{e^{4}\ln(\lambda_{\rm D}/\ell)},$ (101)
where $\ell$, $v_{\rm e}$ and $\lambda_{\rm D}$ are the mean separation of nearest two electrons, i.e., $(n/2)^{-1/3}$, the mean velocity of electrons and the Debye length (see, for example, Appendix B in NMYI paper). By contrast, the period of the Larmor motion in the present situation is given by
$T_{\rm Larmor}=\frac{2\pi r\sin\vartheta}{v_{(\varphi)}}\simeq\frac{2\pi\sqrt{3}r\sin\vartheta}{v_{\rm e}}.$ (102)
The ratio of these two time scales is given by
$\frac{T_{\rm relax}}{T_{\rm Larmor}}=\frac{2\varepsilon_{0}^{2}m_{\rm e}^{2}v_{\rm e}^{4}}{\sqrt{3}\thinspace e^{4}nr\sin\vartheta\ln(\lambda_{\rm D}/\ell)}.$ (103)
By assuming the isotropy of the velocity distribution, we have
$v_{\rm e}^{2}=v_{(r)}^{2}+v_{(\theta)}^{2}+v_{(\varphi)}^{2}\simeq 3v_{(\varphi)}^{2}=\frac{3c^{2}{\cal K}^{2}}{2R^{2}}.$ (104)
From Eq. (51), we have
$\sin\vartheta=\frac{{\cal K}}{\sqrt{{\cal Q}_{\rm M}^{2}+{\cal K}^{2}}}.$ (105)
We find from Fig. 1 that $R_{\rm min}/2<{\cal K}^{2}<R_{\rm max}/2$ holds if we focus on particles bounded within $R_{\rm min}<R<R_{\rm max}$. We are interested in the case of $10<R_{\rm min}<R_{\rm max}<100$, and hence we should assume $5<{\cal K}^{2}<50$. Then, from Eq. (52), we have ${\cal Q}_{\rm M}^{2}\gg{\cal K}^{2}$, and hence we may approximate $\sin\vartheta$ as
$\sin\vartheta\simeq\frac{{\cal K}}{|{\cal Q}_{\rm M}|}.$ (106)
Thus, we have
$\frac{T_{\rm relax}}{T_{\rm Larmor}}=\frac{9\varepsilon_{0}^{2}\left(m_{\rm e}c^{2}\right)^{2}}{2\sqrt{3}\thinspace e^{4}}\times\frac{{\cal K}^{3}|{\cal Q}_{\rm M}|}{r_{\rm g}nR^{5}\ln(\lambda_{\rm D}/\ell)}$ (107)
$\frac{T_{\rm relax}}{T_{\rm Larmor}}$ $=\frac{9\varepsilon_{0}^{2}\left(m_{\rm e}c^{2}\right)^{2}}{2\sqrt{3}\thinspace e^{4}}\times\frac{{\cal K}^{3}|{\cal Q}_{\rm M}|}{r_{\rm g}nR^{5}\ln(\lambda_{\rm D}/\ell)}$
$=\frac{9\varepsilon_{0}^{2}c^{3}m_{\rm e}{\cal K}^{3}}{\sqrt{3}e^{3}nR^{5}\ln(\lambda_{\rm D}/\ell)}\left(\frac{N^{5}}{N-1}\right)^{1\over 2}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)|B_{N}|$ (108)
If $T_{\rm relax}\gg T_{\rm Larmor}$, or equivalently,
$n$ $\ll\frac{9\varepsilon_{0}^{2}c^{3}m_{\rm e}{\cal K}^{3}}{\sqrt{3}e^{3}R^{5}\ln(\lambda_{\rm D}/\ell)}\left(\frac{N^{5}}{N-1}\right)^{1\over 2}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)|B_{N}|$
$\simeq\frac{3.04\times 10^{21}}{\ln(\lambda_{\rm D}/\ell)}\left(\frac{100r_{\rm g}}{r}\right)^{2}\left(\frac{10}{N}\right)^{2}\left(\frac{{\cal K}}{5}\right)^{3}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)\left(\frac{|B_{N}|}{10^{-3}{\rm T}}\right)[{\rm m}^{-3}]$ (109)
holds, the plasma can be regarded to be collisionless, where ,in the last equality, we have assumed, $N\gg 1$. This condition is satisfied in the situations of our interest.
V.3 Magnetic field produced by the electric current caused by charged test particles
The electric current is caused by the motions of charged particles so as to cancel out the background monopole magnetic field. If the electric current is too large, the hovering motions in the plasma lump is highly affected. Thus we consider a lump of charged particles hovering on a black hole and estimate the magnetic field generated by the motion of the charged particles. In order to simplify the calculations, we assume the lump occupies a domain specified by $r_{\rm i}\leq r\leq r_{\rm i}+l$ and $0\leq\theta\leq\Theta$ with
$2r_{\rm i}\sin\Theta=l$ (110)
(see Fig. 2). A particle with $\cal K$} moves on a cone with a half apex angle $\vartheta$ given by Eq. (51). The electric currents generated by charged test particles cancel out inside the lump, and hence the electric current on the surface $\theta=\Theta$ remains.
We assume the system is static and axisymmetric, or equivalently, the four-vector potential $A_{a}$ does not depend on $t$ and $\varphi$, and $j^{\alpha}\propto\delta_{3}^{\alpha}$. Then, Maxwell’s equations are given as
$\sum_{i,j=1}^{3}\frac{\partial}{\partial x^{i}}\left(\sqrt{-g}g^{00}g^{ij}\frac{\partial A_{0}}{\partial x^{j}}\right)$ $=0,$ (111)
$\frac{\partial}{\partial\theta}\left(\sqrt{-g}F^{12}\right)$ $=0,$ (112)
$\frac{\partial}{\partial r}\left(\sqrt{-g}F^{21}\right)$ $=0,$ (113)
$\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}F^{13}\right)+\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\left(\sin\theta F^{23}\right)$ $=-\frac{4\pi}{c^{2}\varepsilon_{0}}j^{3}.$ (114)
If $A_{0}$, $A_{1}$ and $A_{2}$ are constant, the first, second and third equations are trivially satisfied, and we adopt these trivial solutions. The forth equation is rewritten in the form
$\frac{\partial}{\partial r}\left(f(r)\frac{\partial A_{\varphi}}{\partial r}\right)+\frac{\sin\theta}{r^{2}}\frac{\partial}{\partial\theta}\left(\frac{1}{\sin\theta}\frac{\partial A_{\varphi}}{\partial\theta}\right)=-\frac{4\pi}{c^{2}\varepsilon_{0}}j_{\varphi}.$ (115)
We assume the domain of $r\gg r_{\rm g}$, or equivalently, $f(r)\simeq 1$. Then, Eq. (115) becomes Ampere’s law in the flat spacetime. Since our purpose is the order estimate of the magnetic field generated by the motion of charged particles, this assumption will not lead to serious incorrectness.
Then a circular electric current $dI$ going through a domain $r\in[r,r+dr]$, $\theta\in[\Theta-\vartheta,\Theta]$ and $\varphi\in[0,2\pi]$ is given as
$dI=\frac{1}{2}qnv_{(\varphi)}r\vartheta dr,$ (116)
where $n$ is the number density of the charged particles with electric charge $q$, and $v_{(\varphi)}$ is defined as Eq. (60) (also see Fig. 2 and Appendix B).
Figure 2: A clump of plasma hovering on a black hole is depicted, where $\rho:=\sqrt{x^{2}+y^{2}}$. It occupies a domain $r_{\rm i}\leq r\leq r_{\rm i}+l$ and $0\leq\theta\leq\Theta$.
Then the magnetic field generated at the center of this clump, $(r,\theta)=\left(r_{\rm i}+l/2,0\right)$, by the circular electric current is parallel to the $z$-axis and its $z$-component $db^{z}$ is given as
$db^{z}=\frac{dIr^{2}\sin^{2}\Theta}{4c^{2}\varepsilon_{0}}\left[r^{2}\sin^{2}\Theta+\left(r_{\rm i}+\frac{l}{2}-r\cos\Theta\right)^{2}\right]^{-3/2}$ (117)
We need to know the upper bound of the field strength generated by the motion of charged particles in the lump. Since the maximal absolute value of $v_{(\varphi)}$ with the energy fixed is realized in the case that the charged particles moves circularly, we assume the circular motion of the charged particles. In the case of the circular motion, the orbital radius is the minimum of the effective potential $V(R;{\cal K},{\cal Q}_{\rm E}=0)$ defined as Eq. (42), and hence, from Eq. (64), we have
${\cal K}^{2}=\frac{R^{2}}{2R-3}$ (118)
From this equation and Eq. (51), we have
$\vartheta=\arctan\frac{{\cal K}}{|{\cal Q}_{\rm M}|}=\arctan\left(\frac{R}{|{\cal Q}_{\rm M}|\sqrt{2R-3}}\right).$ (119)
From Eqs. (50) and (51), we have
$v_{(\varphi)}=-\frac{c{\cal Q}_{\rm M}}{R}\tan\vartheta=-\frac{c~{\rm sgn}({\cal Q}_{\rm M})}{\sqrt{2R-3}},$ (120)
where sgn(x)(x) is the sign function. By assuming $0<\vartheta\ll 1$ or equivalently, $|{\cal Q}_{\rm M}|\gg R$, since
$\vartheta$ $\simeq\frac{R}{|{\cal Q}_{\rm M}|\sqrt{2R-3}}$ (121)
is a good approximation, we have
$b^{z}=-\frac{nqr_{\rm g}\sin^{2}\Theta}{8c\varepsilon_{0}{\cal Q}_{\rm M}}\int_{R_{\rm i}}^{R_{\rm i}+L}\frac{R^{4}dR}{(2R-3)}\left[R^{2}-2R\left(R_{\rm i}+\dfrac{L}{2}\right)\cos\Theta+\left(R_{\rm i}+\dfrac{L}{2}\right)^{2}\right]^{-3/2},$ (122)
where $R_{\rm i}=r_{\rm i}/r_{\rm g}$ and $L=l/r_{\rm g}$. In order to estimate the integral in R.H.S. of Eq. (122), we assume $L\ll R_{\rm i}$ and then obtain
$b^{z}=-\frac{nqr_{\rm g}}{8c\varepsilon_{0}{\cal Q}_{\rm M}}\times\frac{R_{\rm i}^{2}}{\sqrt{2}\left(2R_{\rm i}-3\right)}\left[1+{\cal O}\left(\frac{L}{R_{\rm i}}\right)\right]\simeq-\frac{nqr_{\rm g}R_{\rm i}}{16\sqrt{2}\thinspace c\varepsilon_{0}{\cal Q}_{\rm M}},$ (123)
where in the last equality, we use the assumption of $R_{\rm i}\gg 1$. Then, by assuming $R_{\rm i}\gg 1$ and $N\gg 1$, we have
$\left|\frac{b^{z}}{B^{r}|_{r=r_{i}}}\right|$ $\simeq\frac{nqr_{\rm g}R_{\rm i}}{16\sqrt{2}\thinspace c\varepsilon_{0}|{\cal Q}_{\rm M}|}\times\frac{R_{\rm i}^{2}}{N^{2}B_{N}}\simeq\frac{nmR_{\rm i}^{3}}{16\sqrt{2}\thinspace\varepsilon_{0}B_{N}^{2}N^{4}},$ (124)
where we have used Eq. (7) in the last equality. If $|b^{z}/B^{r}|_{r=r_{i}}|\ll 1$, or equivalently,
$n$ $\ll\frac{16\sqrt{2}\thinspace\varepsilon_{0}B_{N}^{2}N^{4}}{mR_{\rm i}^{3}}$
$\simeq 1.20\times 10^{9}\left(\frac{m_{\rm p}}{m}\right)\left(\frac{100r_{\rm g}}{r_{\rm i}}\right)^{3}\left(\frac{B_{N}}{10^{-3}{\rm T}}\right)^{2}\left(\frac{N}{10}\right)^{4}[{\rm m}^{-3}]$ (125)
holds, our analysis can be valid. This condition may be satisfied in a typical HII region.
V.4 Backreaction due to the electromagnetic radiation
From the results obtained in Sec. III.3, we can estimate the frequency of the electromagnetic radiation generated by the charged test particle. The frequency $\nu$ of the angular motion of the test particle is estimated to be
$\nu$ $=\frac{1}{2\pi}\left|\frac{d\varphi}{d\tau}\right|=\frac{\left|u_{\varphi}\right|}{2\pi r^{2}\sin^{2}\vartheta}$
$=\frac{c}{2\pi r_{\rm g}}\sqrt{{\cal Q}_{\rm M}^{2}+{\cal K}^{2}}\left(\frac{r_{\rm g}}{r}\right)^{2}$
$\simeq 2.8\times 10^{5}\left(\frac{100r_{\rm g}}{r}\right)^{2}\left(\frac{N}{10}\right)^{2}\left(\frac{q}{e}\right)\left(\frac{m_{\rm e}}{m}\right)\left(\frac{B^{r}|_{r=Nr_{\rm g}}}{10^{-3}{\rm T}}\right){\rm Hz},$ (126)
where we have used Eqs. (48), (49) and (51) in the third equality and Eq. (52) with the replacement of $m_{\rm p}$ by $m_{\rm e}$ in the last equality, and ${\cal Q}_{\rm M}^{2}\gg{\cal K}^{2}$ and $N\gg 1$ have also been assumed in the last equality. From this result, we find that the radiation of very low frequencies may be generated by electrons in this system, if the strength of the magnetic field is the same order as that in the neighborhood of SgrA* or of the center of M87 galaxy.
The rotational velocity of the circular motion is given by
$v_{(\varphi)}^{2}=\frac{c^{2}{\cal K}^{2}}{R^{2}}=\frac{c^{2}}{2R-3},$ (127)
where we have used Eq. (118) in the last equality. Hence, the rotational velocity is much less than the speed of light if the radial position of the test particle satisfies $r\gg r_{\rm g}$.
In the case that the motion of the charged test particle is non-relativistic, the radiation power $I_{\rm EM}$ of the magnetic bremsstrahlung by one particle is estimated by the dipole formula
$I_{\rm EM}=\frac{1}{6\pi c^{3}\varepsilon_{0}}\sum_{i=1}^{3}\frac{d^{2}D^{i}}{dt^{2}}\frac{d^{2}D^{i}}{dt^{2}},$ (128)
where $D^{i}$ is the electric dipole moment defined as
$D^{i}=qx^{i}$ (129)
with the spatial trajectory of the charged test particle, $x^{i}=x^{i}(t)$, which is approximately given as
$x^{1}$ $=r\sin\vartheta\cos\left(2\pi\nu t\right),$ (130)
$x^{2}$ $=r\sin\vartheta\sin\left(2\pi\nu t\right),$ (131)
$x^{3}$ $=r\cos\vartheta.$ (132)
By assuming $\left|\dfrac{dr}{dt}\right|\ll\nu r$, we have
$I_{\rm EM}=\frac{(2\pi\nu)^{4}q^{2}r^{2}\sin^{2}\vartheta}{6\pi c^{3}\varepsilon_{0}}.$ (133)
Then, by using Eq. (126), we have
$I_{\rm EM}$ $=\frac{ce^{2}}{6\pi\varepsilon_{0}r_{\rm g}^{2}}\frac{{\cal K}^{2}({\cal Q}_{\rm M}^{2}+{\cal K}^{2})}{R^{6}}\left(\frac{q}{e}\right)^{2}$
$\simeq\frac{e^{4}B_{N}^{2}}{12\pi\varepsilon_{0}cm_{\rm e}^{2}}\frac{N^{4}}{R^{5}}\left(\frac{m_{\rm e}}{m}\right)^{2}\left(\frac{q}{e}\right)^{4}$
$\simeq 4.95\times 10^{-8}\left(\frac{N}{10}\right)^{4}\left(\frac{100r_{\rm g}}{r}\right)^{5}\left(\frac{m_{\rm e}}{m}\right)^{2}\left(\frac{q}{e}\right)^{4}\left(\frac{B_{N}}{10^{-3}{\rm T}}\right)^{2}{\rm eV/s},$ (134)
[…]
VI Summary and discussions
We studied the motion of a charged test particle around a spherically symmetric black hole in a monopole magnetic field. Its radial motion is completely the same as that of the charged test particle in the case with no magnetic field. This fact immediately leads to the same results on the electrification of a black hole as the case with no magnetic field. If plasma composed of non-relativistic protons and electrons whose temperatures are not so different from each other surrounds a spherically symmetric black hole in the monopole magnetic field, selective accretions of protons will occur due to the difference of the masses between a proton and an electron. Since $T_{\rm e}<T_{\rm p}<5T_{\rm e}$ will hold in the situation of our interestMos ; Dexter ; Yuan-Narayan , the black hole will acquire positive charge. The total amount of charge will be very small amount but its effect on the charged particles is very large. For example, the radius of ISCO of an electron can be one order larger than those of the electrically neutral test particlepaper-I . We need to take into account these effects on the plasma surrounding an electrified black hole.
By contrast to the radial motion, the motion in angular directions of the charged test particles in the monopole magnetic field is very different from that in the case with no magnetic field. The motion of an electrically neutral test particle is restricted to an equatorial plane, whereas that of a charged test particle is restricted on a cone whose half apex angle $\vartheta$ is determined by the magnetic charge of the black hole, the electric charge and the angular momentum of the particle. In the situation with the magnetic field whose strength is similar to SgrA* or the center of M87 galaxiy, the apex angle $\vartheta$ of the cone is much less than one radian, if the specific angular momentum of the charged particle is much less than $10^{6}cr_{\rm g}$. Here note that the orbital radial position $r$ of a circularly moving particle with the specific angular momentum $K$ is given as $r\simeq K^{2}/cr_{\rm g}$. Hence the particle bounded within the domain of $r<10^{4}r_{\rm g}$ will have a specific angular momentum less than $10^{2}cr_{\rm g}$. These results imply that a charged test particle of a circular motion at the radial position $r$ does not stay on an equatorial plane but hovers on a black hole in a very narrow region of the tangential extent, or in other words, the Larmor radius, $r\vartheta$, which is much less than $r$. The distribution of plasma around the black hole in the monopole magnetic field may be rather different from the case with no magnetic field.
By the same reason, a plasma lump can hover on the black hole in the monopole magnetic field, if the size of the lump is much larger than the Larmor radii of the constituent charged particles $r\vartheta$. Here note that a plasma particle of the Larmor radii $r\vartheta$ has kinetic energy equal to that of the Keplerian motion of the orbital radius $r$. Due to this fact, the mean kinetic energy of plasma particles, or equivalently, the local temperature of the plasma lump may be extremely high. For example, the temperature of protons may be the order of $10^{10}$K when the radial position of the plasma lump is equal to about $100r_{\rm g}$ in the case of the black hole with the mass same as that of SgrA*. Although the temperature of the plasma lump is very high, the thermal radiation will not be emitted since the plasma is too diffuse to be in thermal equilibrium with photons in the situation we consider. It seems to be difficult to observe this high temperature of the plasma, even if it is the case.
It is a crucial prediction that the average velocity of electrons will be equal to that of protons, or in other words, the temperature of electrons is much lower than that of protons, in a plasma lump. This situation is very different from the assumed situation in the discussion of the electrification of a static black hole, where we have assumed that the velocity distribution of electrons and that of protons are Maxwell one with not so different temperatures from each other. Hence, we should note that bounded lumps of collisionless plasma will not be suppliers of positive electric charge to the black hole. However, as shown in Sec. V.4, electrons in a plasma lump can be selectively accreted on the black hole through the energy loss due to the electromagnetic radiation. Thus, lumps of collisionless plasma bounded around a black hole may be suppliers of negative electric charges to the black hole in a monopole magnetic field.
Acknowledgments
We are grateful to colleagues in the astrophysics and gravity group in Osaka Metropolitan University. This work was supported by JSPS KAKENHI Grants No. JP21K03557 (K.N.), No. JP21H05189 (H.Y.), No. JP22H01220 (H.Y.), and MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165 (K.N. and H.Y.).
Appendix A About a test particle initially moving outward
We consider a test particle initially moving outward. It should be bounded, or equivalently, ${\cal E}^{2}<1$ holds so that it falls into the black hole. Equation (66) implies that the bounded test particle should satisfy
$\frac{v_{(r)}^{2}}{c^{2}}<\frac{1}{1-1/R}-1-\frac{{\cal K}^{2}}{R^{2}},$ (135)
or equivalently
${\cal K}^{2}<R^{2}\left(\frac{1}{R-1}-\frac{v_{(r)}^{2}}{c^{2}}\right).$ (136)
In Fig. 3, we depict the parameter region in which both Eq. (76) and Eq. (136) are satisfied as a shaded domain in the $(v_{(r)}^{2}$${\cal K}^{2})$-plane, where $v_{(r)}^{2}=V^{2}$ is a root of
${\cal K}_{\rm max}^{2}(v_{(r)}^{2},R)=R^{2}\left(\frac{1}{R-1}-\frac{v_{(r)}^{2}}{c^{2}}\right).$ (137)
For $R\gg 1$, by using Eq. (77), we have
$V^{2}=\frac{c^{2}}{R}\left[1-\frac{4}{R}+{\cal O}\left(R^{-2}\right)\right].$ (138)
Figure 3: The parameter region in which both Eq. (76) and Eq. (136) are satisfied as a shaded domain in the $(v_{(r)}^{2}$${\cal K}^{2})$-plane.
Thus, in the case of $R\gg 1$, the probability $P_{\rm out}$ of the test particle initially moving outward falling into the black hole is given as
$P_{\rm out}$ $=\int_{0}^{V}dv_{(r)}\left(\frac{m}{2\pi k_{\rm B}T}\right)^{1\over 2}\exp\left(-\frac{m}{2k_{\rm B}T}v_{(r)}^{2}\right)$
$~~\times\int_{0}^{2\pi}d\phi\int_{0}^{{\cal K}_{\rm max}^{2}\left(v_{(r)}^{2},R\right)}d{\cal K}^{2}\frac{mc^{2}}{4\pi k_{\rm B}TR^{2}}\exp\left(-\frac{mc^{2}}{2k_{\rm B}TR^{2}}{\cal K}^{2}\right)$
$+\int_{V}^{\frac{c}{\sqrt{R-1}}}dv_{(r)}\left(\frac{m}{2\pi k_{\rm B}T}\right)^{1\over 2}\exp\left(-\frac{m}{2k_{\rm B}T}v_{(r)}^{2}\right)$
$~~\times\int_{0}^{2\pi}d\phi\int_{0}^{\frac{R^{2}}{R-1}-\frac{R^{2}v_{(r)}^{2}}{c^{2}}}d{\cal K}^{2}\frac{mc^{2}}{4\pi k_{\rm B}TR^{2}}\exp\left(-\frac{mc^{2}}{2k_{\rm B}TR^{2}}{\cal K}^{2}\right)$
$\simeq\sqrt{\frac{2}{\pi}}\left(\frac{mc^{2}}{k_{\rm B}TR}\right)^{1/2}\times\frac{mc^{2}}{k_{\rm B}TR^{2}}\left[1+{\cal O}\left(R^{-1}\right)\right].$ (139)
If $R\gg mc^{2}/k_{\rm B}T$ holds, this probability is much less than that of the test particle initially moving inward falling into the black hole.
Appendix B Electric current on the surface of a clump
Figure 4: A neighborhood of the surface at $x=0$ of a clump which occupies the domain $x<0$ is depicted. The centers of the circular orbits of particles are located in the domain $x\leq-l$. The velocities of particles located at $x$ are isotropic in the case of $x\leq-2l$, whereas those are anisotropic in the case of $2l<x\leq 0$.
In the case that charged test particles distribute homogeneously, the electric current exist only on the surface of the clump considered in Sec. V.3. We focus on the electric charge $q$ of test particles are identical to each other and positive, and hence the orbital radii of the charged particles are identical to $l:=r\vartheta$. Then we consider the situation depicted in Fig. 4.
The velocities of particles in 2l<x<0-2l<x<0 are given as
$\bm{v}=\left(-v_{(\varphi)}\sin\psi,-v_{(\varphi)}\cos\psi\right)_{xy}~~~~{\rm with}~~\psi_{-}<\psi<\psi_{+}$ (140)
where
$\psi_{-}$ $=\frac{1}{2}\pi+\arcsin\left(\frac{x+l}{l}\right),$ (141)
$\psi_{+}$ $=\frac{3}{2}\pi-\arcsin\left(\frac{x+l}{l}\right).$ (142)
Hence, the average velocity of the particle at $x$ is given as
$\langle\bm{v}\rangle$ $=-\frac{v_{(\varphi)}}{2\pi}\left(\int_{\psi_{-}}^{\psi_{+}}\sin\psi d\psi,\int_{\psi_{-}}^{\psi_{+}}\cos\psi d\psi\right)_{xy}$
$=\frac{v_{(\varphi)}}{2\pi}\left(0,+2\cos\left[\arcsin\left(\frac{x+l}{l}\right)\right]\right)_{xy}$ (143)
Then the electric current for $(r,r+dr)$ is equal to
$dI=2qndr\int_{-2l}^{0}\frac{v_{(\varphi)}}{2\pi}\cos\left[\arcsin\left(\frac{x+l}{l}\right)\right]dx=\frac{1}{2}qnv_{(\varphi)}ldr=\frac{1}{2}qnv_{(\varphi)}r\vartheta dr,$ (144)
where $n$ is the number density of the charged particle of $q$.